### v1

# Visual Writing: Writing by Manipulating Visual Representations of Stories

Abstract.
We introduce “visual writing”, an approach to writing stories by manipulating visuals instead of words. Visual writing relies on editable visual representations of time, entities, events, and locations to offer representations more suited to specific editing tasks. We propose a taxonomy for these representations and implement a prototype software supporting the visual writing workflow. The system allows writers to edit the story by alternating between modifying the text and manipulating visual representations to edit entities, actions, locations, and order of events. We evaluate this workflow with eight creative writers and find visual writing can help find specific passages, keep track of story elements, specify edits, and explore story variations in a way that encourages creativity.
1. Introduction
Writing stories and exploring alternative narratives can be challenging. Although using large language models (LLMs) can help writers (Chakrabarty et al., 2024), it remains difficult for users to formulate their intent using words (Dang et al., 2022b) and review the generated text to ensure it aligns with their intent (Masson et al., 2024).
Consider a task as simple as changing the location of a character in a story: the cat no longer goes to the barn but instead wanders about the lake. Using “replace all” instances of ‘barn’ is insufficient because substituting a location comes with implications. For example, descriptions and actions pertaining to the barn would need to be updated to reflect the new context of being by the lake. Further, not all instances necessarily need to be impacted. The fact that the cat no longer visits the barn should not result in the barn being completely removed from the story. Instead, writers must carefully plan to find all and only the references to the location associated with the character’s change of action and make tedious revisions to the text to maintain spatial and semantic consistencies. This busywork is prone to errors, and writers often resort to creating external documents such as maps, spreadsheets, and timelines that they painstakingly update to keep track of the characters, events, and locations (Marshall, 2014; Ackerman, 2017; Neuwirth and Kaufer, 1989; SaveTheCat, 2014). While LLMs could help with this busywork, expressing distinct changes can be challenging with prompts (Masson et al., 2024), and the extent and suitability of the changes made by the generative model could be difficult to verify.
The challenge stems from the mismatch between the reasoning and representation spaces: in our above example, the writer performs spatial reasoning using a textual representation. A spatial representation, such as a map of the story world, would be more suited for such a task, where changing the location of the character becomes as simple as moving its position on the map. In text, edits relevant to space and locations in the story are comparatively challenging to accomplish, which can be a deterrent for writers wanting to change the narrative and explore variations. And this same argument applies to all other aspects of a story such as time, events, and characters, for which we posit superior representations can be designed.
We define visual writing as the use of visual representations of story elements as an input medium for expressing writing intents. For each story element, we explore the use of a visual representation that conceptually aligns with that element. Depending on their editing task, users can decide to edit the text of the story, in which case the visual representations are updated to match the new text. Alternatively, they can manipulate the visual representations through direct manipulation to suggest changes to the text and rapidly explore variations. We discuss a taxonomy of these visual representations, drawing on the story elements framework from narratology (§3). In a prototype software (§4), we implement three of these visual representations: (1) a diagrammatic view of the entities in the story and their relationships, allowing to add, remove, and edit characters and add new actions between them (Figure 1.c); (2) a spatial view of the locations in the story and the locations of the entities, allowing to add new locations and move entities (Figure 1.d); and (3) a timeline of the events in the story, allowing to reorder the events, quickly find specific scenes, and precisely modify the entities and locations of the selected scenes by leveraging the other two views (Figure 1.b).
We validate this approach in a user study with eight creative writers (§5). Results show participants appreciated the visual representations to locate scenes and keep track of story elements while writing. They were successful in expressing their editing intents through direct manipulations of the visual elements. Furthermore, participants appreciated this approach to edit text from an exploration and creativity perspective. This points to the potential of visual representations for writing and calls for more work in that direction.
2. Related Work
Visual writing draws inspiration from work on visualizing stories and efforts to give alternative representations to edit content. Below, we review the literature on these topics.
2.1. Visualizing Stories
Visual representations of stories and story patterns have a long-standing presence in writing, with staple examples including Freytag’s pyramid story (Freytag, 1895) and abstract diagrams capturing the “shape of the story” (Pope, 1998; Comberg, [n. d.]; Reagan et al., 2016). Such visual representations are not only useful for teaching the craft of storytelling, they are also considered as useful aids to support the writing process. Referring to the visual representations, screenplay writer E. Williams says: “I find story shape to be a great tool […] to make sure my story holds together” (Williams, [n. d.]).
There exist many techniques to visualize stories, spanning from classical charts like word clouds, to more artistic visualizations like calligrams (Brath, 2021). Often, these visualizations summarize entire stories in an attempt to highlight a specific facet (e.g., emotions (Maharjan et al., 2018) or topic recurrence (Zhu, 2013)). For example, Storyline visualizations (Munroe, 2009; Tanahashi and Kwan-Liu Ma, 2012) depict the evolution of the relationships between characters through timelines. StoryCake (Qiang and Bingjie, 2016) uses a hierarchical plot to visualize the structure of nonlinear stories. Story Curves (Kim et al., 2018) visualizes non-linear narratives by showing both the order in which the events are told and the order of the underlying story. StoryPrint (Watson et al., 2019) shows the scenes and character emotions through circular timelines. Each of these visual representations presents the ability to support readers, writers, and analysts alike in appreciating or even analyzing the stories from a specific lens.
Other story visualizations have been used in tools to help during script and story writing. For example, CARDINAL (Marti et al., 2018) is a tool to help screenwriters visualize their script through different views, including a visualization of the interactions between characters and an animated 2D and 3D views to play the scene and visualize locations and movements of the characters. Similarly, Portrayal (Hoque et al., 2023) attempts to help writers develop their characters by extracting and presenting information related to the characters in interactive visualizations.
Our work explores the potential of expanding the role of such visualizations to also be manipulable. That is, by mapping visualization elements directly to entities, events, and locations in the text story, we explore using visualizations as the medium to modify the story. This approach is motivated by the potential of visual representations to express intent, which we discuss next.
2.2. Editing Content Through Alternative Representations
Programming might be the domain that most embraced the use of alternative representations. Approaches include visual programming (Myers, 1986; Burnett and McIntyre, 1995), which allows generating programs by manipulating elements in more than one dimension, code projections (Gobert and Beaudouin-Lafon, 2023; Ko and Myers, 2006) which augment programming languages with tools and widgets that are alternative representations of fragments of code, and bi-directional programming (Mayer et al., 2018; Hempel et al., 2019) where the output of a program can be manipulated to edit the program itself. Similar to our work, the motivation behind these approaches is that some tasks are easier done using different representations: it is easier to find a colour using a colour picker than by typing its hexadecimal code, and it is easier to move an object on the screen through direct manipulation (Shneiderman, 1983) than by typing its coordinates.
In the creative writing domain, this idea of manipulating projections or visual representations to edit stories is largely untapped, which is surprising given the reliance on visual representations to capture story concepts (Freytag, 1895; Comberg, [n. d.]). Instead, most writing support tools ask users to manipulate text using text, typically through prompts or suggestions (Calderwood et al., 2020; Buschek et al., 2021; Yuan et al., 2022; Dang et al., 2023; Lee et al., 2024; Reza et al., 2024). There are some exceptions, such as the proposal of Kempen et al. (1987) to represent and manipulate text in different ways, such as manipulating the syntax tree to allow restructuring sentences. Similarly, Arnold et al. (2021) suggested that words could be directly manipulated to be reordered. Beyond syntactic level modifications, Dang et al. (2022a) proposed a system that allows reorganizing, merging, and removing paragraphs by manipulating textual summaries instead of the text. In the same vein, VISAR (Zhang et al., 2023) automatically generates an interactive visual outline of an argumentative text, which can be edited to change the structure of the argument. Finally, TaleBrush (Chung et al., 2022) leverages a canvas where sketching allows modulating the fortune of characters in a story.
Our work contributes to this line of research by formalizing the concept of visual writing. The concept is rooted in the same principles as visual programming and code projections, positing that elements of a story become easier to understand and manipulate when presented in alternative forms. Specifically, we design visualizations grounded in narratology to explore this approach.
3. Taxonomy of Visual Representations
Narratology separates the chronological sequence of the story (fabula) from the narrative sequence (syuzhet) of the story. At its core, a fabula consists of four main story elements: events, actors111In narratology, actors refer to “agents that perform actions. They are not necessarily human” (Bal, 1997). To avoid confusion, we use the term entity instead throughout this paper., time, and location (Bal, 1997). Similarly, we define visual writing as the use of visuals that represent and manipulate these four elements.
In their simplest form, visual representations can depict elements one at a time. For example, a list of locations would allow adding, removing, and editing the characteristics of each location. And a list of entities, similar to the character sheets writers often create to plan their stories (SaveTheCat, 2014), would allow adding and removing entities. However, many important aspects of stories are encoded through the interaction between elements (e.g., the time of an event or the location of an entity). Below, using the narratology lens, we discuss a non-exhaustive taxonomy of possible visual representations as an initial guide to inform the design of tools supporting visual writing.
Spatial representations depict elements based on their location. For example, a spatial representation could be a map with dots, where each dot represents where events happened and where entities are located. Or it could be more abstract and only represent distances between elements. Spatial representations are used by systems like CARDINAL (Marti et al., 2018) to show the relative positioning of characters within a scene.
Event-driven representations depict elements based on the events they relate to. For example, a node graph could be an event-driven representation if each node represents an event and the edges encode another element such as the entities (in which case, connected nodes would represent events affecting the same entity) or time (in which case the directed edges would indicate the flow of the events). Some examples of event-driven representations are lists of outline points and event graphs.
Entity-driven representations depict elements based on the entities they relate to. For example, a node graph could be an entity-driven representation if each node represents an entity and the edges encode another element such as the events (in which case, each edge is an event connecting the entities involved in the event) or time (in which case, the directed edges indicate the order in which the entities intervene in the story). Entity-driven representations, such as character sheets and family trees, are often created by writers to keep track of character details (SaveTheCat, 2014).
Temporal representations depict elements based on the time they happen in the fictional world of the story or in the order they are presented in the text. For example, a timeline could arrange elements such as events (in which case, the events would be positioned based on when they occur) or entities (in which case, entities would be positioned based on when they appear or when they perform an action). Temporal representations have been more commonly used by previous work on visualizing stories such as Storylines (Munroe, 2009), StoryCake (Qiang and Bingjie, 2016), and StoryPrint (Watson et al., 2019).
The design space for the above categories is vast, and there exist many visualizations to bear inspiration from. Our goal is not to argue which specific design is best—we leave it to future work to explore different variations within these categories (and beyond). Instead, our work focuses on the potential of visual writing, i.e. the use of alternative representation of stories, as an input medium for expressing intent.
4. The Visual Writing Prototype
To test the potential of visual writing, we developed a prototype system, shown in Figure 1, implementing three visual representations: a diagrammatic view of the entities and actions (Entity–Event), a spatial view of the locations and entities (Location–Entity), and a timeline of the events (Time–Event). The focus of the prototype was to help writers explore variations and perform spatial and temporal edits via visual writing. As such, the system allowed specifying edits through direct manipulations of the graphical elements (Shneiderman, 1982; Hutchins et al., 1985) and followed the principles to design creativity support tools (Shneiderman et al., 2006). Below, we detail the design and features of the system. All the examples and screenshots were taken using the system on the first three paragraphs of the story Alice’s Adventures in Wonderland by Lewis Carroll.
4.1. Entities and Actions View
This event-driven representation shows all the entities in the story as nodes with their name and an emoji representing them (Figure 1.c). The actions between the entities are represented as directed edges with a label describing the action in one or two words. By default, the nodes are arranged using a force layout to prevent overlaps, but users can rearrange the position of the entities by dragging them.
4.1.1. Editing the Traits of an Entity
Selecting an entity expands its node to reveal traits related to the entity similar to a list of personality traits which would be present on a character sheet (e.g., “curious”, “adventurous”, “chatty”). The intensity of each trait is rated from 1 to 10. This intensity can be changed by moving the slider beside the trait. This results in editing the text of the story to reflect the new intensity of the trait (Figure 2.a).
4.1.2. Adding and Removing an Entity
Adding an entity is done by double-clicking anywhere on the canvas which opens a text input box to indicate the name of the new entity. Once the name is entered, the entity is created at the location of the click. Removing an entity is done by selecting the entity and then pressing “delete” or “backspace” on the keyboard. This results in updating the text of the story so that the entity is effectively removed (Figure 2.b).
4.1.3. Adding and Removing Actions between Entities
Adding an action is done by creating an edge between two nodes. The direction of the edge indicates the source and target of the action. An action can also be created for the same character by connecting a node to itself. This helps represent actions done by an entity with no target (e.g., walking or thinking). An action can be edited at all times by double-clicking its label and changing its description. Similarly, an action can be removed by being selected and then pressing “delete” or “backspace”. As soon as the edge is created, edited, or deleted, the story text is updated to reflect the change (Figure 2.c).
4.2. Locations and Entities View
The location view is a spatial representation accessed through a tab (Figure 1.d). It shows all the locations in the story as nodes with their name and an emoji representing them. Similar to the entities and actions view (Section 4.1), the view shows all entities as smaller nodes, but their position is conditioned by their location in the story. Entities that go to different locations in the story are duplicated to be represented in all of their locations.
4.2.1. Moving an Entity
Entities can be dragged and moved around the view. If the entity is released on top of a location, then that entity is moved to that location (Figure 3.b-c). Otherwise, the entity goes back to its original location. Once the entity is moved, the story text is updated to reflect the change in location.
4.2.2. Adding a Location
A new location can be added by double-clicking anywhere on the canvas (Figure 3.a). This opens a text input field to specify the name of the location. Once the location is created, entities can be moved to it.
4.3. Timeline of Events View
The timeline view is a temporal representation showing the events in the story represented as vertical lines with the emojis of the entities involved in the event on either side of the line (Figure 1.b). These lines are organized one after the other based on the order in which they are presented in the text. The timeline view is always visible. It allows selecting one or multiple events, in which case the other views are updated to fade out the entities and locations not involved in the selected events. Selecting events also forces future modifications to impact only the selected events.
4.3.1. Reordering Events
Events can be dragged to be reorganized temporally. When an event is dragged, it slides horizontally and can be placed between other events (Figure 4.b-c). Once the event is moved, the story text is updated so that the event happens when indicated in the timeline.
4.3.2. Selecting Events
The timeline also allows finding specific scenes and selecting them to allow precise editing. Hovering over events highlights the corresponding sentence in the story (Figure 4.a) and the corresponding entities and locations in the event and location views. Clicking selects the event, and dragging selects multiple events. Once one or multiple events are selected, any visual edits made in the views will target these selected events. For example, selecting an event and then moving an entity in the location view will update the story text so that the entity moved exactly during the selected event.
4.4. Bi-directional Editor
The rest of the interface is composed of a text editor on the left, a history tree at the bottom to navigate different versions of the story, and interface buttons at the centre to refresh the visuals or rewrite the story (Figure 1.a).
4.4.1. Highlighting Visual Elements and Sentences on Hover
Similar to how hovering over the events in the timeline highlights the sentences, hovering over the sentences in the text editor highlights the corresponding events, entities, and locations in the different views. This allows for quickly finding different scenes. Similarly, placing the caret in one sentence will act as if the events in that sentence were selected in the timeline. As such, future edits to the visual representations will be restricted to that sentence.
4.4.2. Updating the Visual Representations from the Text
When the text is updated manually, the visual representations might become desynchronized, showing the story as it was before the modification. In this case, the button to refresh the visuals becomes blue to indicate that the visual model may need to be updated. Clicking it results in re-extracting the information from the text and refreshing the views.
4.4.3. Writing the Story from the Visual Representations
By clicking the “refresh from visuals” button, the story is rewritten from scratch using only the visual representations as reference. As such, the story preserves the same events, locations, and entities, but everything else is rewritten. This allows specifying the skeleton of a story and quickly exploring different ways of phrasing the story.
4.4.4. Track changes
By default, all changes made by editing the visual representations are tracked, such that removed text is struck in red while additions are shown in blue. This allows writers to easily locate and review the changes before accepting (or rejecting) them.
4.4.5. History Tree
To make edits easily reversible and facilitate the exploration of alternatives, the system also implements a history tree capable of storing different versions of the story and returning to them at any time. Every time a modification happens, be it by editing the visual representation or editing the text of the story, a snapshot of the story and visual representation are added to the tree. If the writer selects a previous snapshot, then the story is changed to that snapshot. If, after selecting a previous snapshot, a modification is made to the story, then this creates a new branch to the story. Then, writers can navigate between the branches and different versions by clicking on the nodes of the history tree.
4.5. Implementation
We implemented the visual writing system using TypeScript and React (React, 2013) for the interface, NextUI (NextUI, 2021) for the graphical components, Slate.js (SlateJS, 2016) for the text editor, and the OpenAI library for prompting large language models (OpenAI, 2023). Under the hood, modifications to the visual representations are turned into engineered prompts sent to OpenAI’s “GPT4-o” model.
4.5.1. Extracting Entities, Locations, and Events
The visual writing approach requires a robust, exhaustive, and fast extraction of the information needed for the visual representations. In our tests, a naive approach consisting of giving a large language model the whole story and asking to extract the necessary information failed at all three of these requirements. Instead, we force the model to output JSON with the proper schema to make the extraction robust. We check that the schema is correct and prompt the model again if necessary. To make the extraction exhaustive, we proceed in multiple steps by first asking the model to extract the entities and locations in the story. We then ask the model to extract events between these entities and at those locations. Finally, and most importantly, we split the whole story into sentences to make the extraction fast. We then prompt the model with the story, but we ask it to extract only the information within the sentence. This allows us to execute the prompt requests in parallel. As such, a text with 30 sentences will send 30 different smaller prompts to the model. This also gives us an indication of the position of the extracted information, as we know that it comes from the sentence being extracted.
4.5.2. Engineered Prompts to Edit the Story
The system relies on many engineered prompts such as splitting requests to make them faster, giving the model only a subset of the text to force localized edits, and giving the model the previous abstract representation of the story and the new one to help it understand the edits. We encourage interested readers to review the source code for more details.
5. User Study
We evaluated the visual writing workflow by conducting a user study using our prototype system. Specifically, the study tried to answer the following questions: Are the visual representations understood and helpful? Can people express their editing intents using visual writing? What is the impact of visual writing on exploration and creativity?
Our study is separated into two parts. The first part isolates the use of the three different views in a controlled setting. Because our approach is compatible with other ways to edit text (e.g., no assistance or LLMs) and our main interest is to measure the usefulness of visual writing, we decided against comparing our approach to a baseline. Instead, we collected qualitative insights through semi-structured interviews and explicitly asked participants to compare visual writing with their own workflow. Representations were also evaluated quantitatively using simple usability questionnaires.
The second part of our study was a free-form creative writing task to let participants appropriate the tool and its features.
5.1. Participants
We recruited 8 creative writers from our social network and word of mouth (age range: 20 to 31, M=24; 7 self-identified as female and 1 as male). All had several years of experience with creative writing, including personal writing (N=7), fiction (N=4), scriptwriting (N=2), and poetry (N=2). On a scale from 1-never to 5-always, they reported how often they use visuals (e.g., mind maps, timelines) during their writing process (Mdn=2), how often they find visuals to help during the writing process (Mdn=3), and how frequently they use AI services to help during their writing activities (Mdn=1). Participants reported using a range of types of visuals to help their writing, including mood boards (N=4), timelines (N=3), mind maps (N=2), and reference imagery (N=2).
5.2. Apparatus
Participants joined the study remotely via the Zoom conference software. They accessed the experiment website using their own computer. The sessions were audio and screen recorded, and interactions with the experiment website were also recorded (e.g., features used, commands executed, and answers to questions). The sessions were about 45 minutes long. In appreciation for their time, participants received a CAD$15 gift card. The study was approved by our institutional ethics board.
5.3. Design and Procedure
Introduction (~2min)
After being greeted, the participant was asked to complete the consent form and demographics questionnaire. They were then told that they would use a system to help them edit stories.
Video Tutorial ((1 + 3 tasks) 30 seconds)
Before each task, the participant watched a short video demonstrating the features they were about to use in the task. The video used a different text and different tasks to make sure the participant could not just reproduce what happened in the tutorial. Before the first task, the participant watched an additional video presenting the concept of the tool and the organization of the interface.
Story Editing Tasks (3 tasks 5min)
Immediately after watching the video tutorial, the participant were prompted to complete the task. We had three task blocks for each of the visual representation views: (1) entities and actions view tasks required removing an entity and then adding a new entity and an action with the main character; (2) location and entities view tasks required moving an entity to an existing location and then creating a new location and move another entity to it; (3) timeline view tasks required the reordering of events. The instructions for each task were phrased as questions and in a way that forced participants to engage with the story. For example, one of the tasks related to the timeline read “What if Jack loses his hat when he flies into the sky?”, forcing participants to understand enough about the story to know where those events occur and move the action to the right location. The tasks were counter-balanced across participants.
Post-Task Questionnaire (3 tasks 2min)
After each task, the participant first indicated how successful they felt in accomplishing the task on a 5-point semantic differential scale. Then, they answered five questions on a 5-point scale related to usability, understandability, and integration into their workflow.
Free-form Exploration (5-15min)
After accomplishing all the tasks and having a good grasp of the tool, the participant engaged in a free-form exploration. We provided the beginning of a story and asked the participant to continue the story with no instructions other than to “explore what could happen next”.
Post-Study Questionnaire (2min)
After completing the free-form part of the study, the participant answered one last questionnaire, which included the creativity support index (CSI (Cherry and Latulipe, 2014)) and the raw NASA TLX.
Semi-structured Interview (10min)
At the end, the participant engaged in a semi-structured interview with the experimenter. After giving their immediate impressions of the tool, they were asked about their experience with respect to the main categories of the CSI, like enjoyment, exploration, expressiveness, immersion, and satisfaction given the effort. They were also asked about each representation and their perceived overall advantages and disadvantages of the system.
5.4. Results
Participants completed the task and the free-form exploration without major difficulties. Below, we summarize the quantitative findings using descriptive statistics. Following previous recommendations for study designs similar to ours (Zhu and Kolassa, 2018; Masson et al., 2023), confidence intervals are calculated using the studentized bootstrapping method using 10,000 repetitions.
Success rate
On a 5-point semantic differential scale from 1-unsuccessful to 5-successful, participants rated their success M=4.25 (95% CI: [3.8, 4.6]) at accomplishing the tasks. Specifically, they were most successful at the task about adding and removing entities (M=4.5, 95% CI: [3.8, 5]), followed by changing the locations of entities (M=4.25, 95% CI: [3.4, 4.8]), and reordering events (M=4, 95% CI: [2.8, 4.7]).
Creativity Support Index (CSI)
Participants generated an average CSI score of 71.5 (SD=21.12) for the system and the tasks. Table 1 shows the breakdown of the CSI results. Enjoyment was both the most highly rated factor and also the factor considered most important by participants. Immersion was ranked second most important yet received the lowest score. Possible reasons for these ratings are discussed in Section 6.
| Scale | Avg. Score (SD) | Avg. Counts (SD) |
| Enjoyment | 15.73 (1.84) | 3.50 (1.00) |
| Immersion | 12.18 (3.04) | 3.13 (1.76) |
| Results Worth Effort | 15.28 (2.26) | 3.00 (1.50) |
| Exploration | 13.48 (2.91) | 2.88 (1.27) |
| Expressiveness | 13.70 (2.86) | 2.50 (0.50) |
| Overall CSI Score | 71.50 (21.12) |
Workload (Raw NASA-TLX)
Participants reported high performance (M=7.49 SD=1.95) for low mental demand (M=4.11, SD=2.38), effort (M=3.39, SD=2.17), frustration (M=2.5, SD=1.2), temporal demand (M=1.18, SD=1.04), and physical demand (M=1.18, SD=1.04).
Questionnaire
Overall, answers to the 5-point scale statements from 1-strongly disagree to 5-strongly agree were mostly positive, see Figure 5 for a breakdown of the results. Participants found the timeline the easiest view to understand (Mdn=4) and use (Mdn=4). Participants mostly agreed that the views helped accomplish the tasks (Mdn=5), that they would be useful within their writing process (Mdn=4), and that they would easily be integrated into their writing process (Mdn=4).
6. Discussion
We first discuss our results in light of participants’ comments before opening the discussion to other themes stemming from our observations and discussions with participants.
Visual writing can help find scenes and keep track of events, entities, and locations
Most participants agreed that “The visual representation was easy to understand” for all three views. Particularly, they mentioned how the views helped them. P5 appreciated that you can “click on a person, click on a location, and then see stuff pop up” instead of “having to look through […] many paragraphs […] or doing keyword search”. P4 explained “Just having a visual representation helps, you know, just with the different perspectives”. P3 added “Moving around the timeline was a good way of, kind of figuring out different events”. P1 and P5 also mentioned how they struggle to keep track of everything while writing their stories and that the system would help them keep track of that. For instance, P5 details that currently they “have booklets everywhere, and [they] have it annotated weird […] it’s a mess and this is so much easier. You got like, you know, your locations, your people […] keeping track of things. Organizing who’s where, what’s going on, being able to scroll through, like highlighting part of the text shown on the timeline […], making it easy to understand”.
Visual writing can help specify temporal, spatial, and entity-related edits
Participants reported high success rates when accomplishing the tasks and provided high ratings for statements such as “The visual representation helped me accomplish the tasks” and “The visual representation was easy to use”. This suggests that visual writing facilitates temporal, spatial, and entity-related edits through the views provided in our system. Participant’s comments further support these results. For example, P2 mentioned “If I add in an extra character, it gives you just enough text to build that story with that new character”. P4 appreciated that it made the story consistent when editing locations: “When I was dealing with the change where the hat was supposed to be left at home, the fact that it also updated the sentence where the hat flew off his head that is extremely useful because that’s the sort of inconsistencies that can easily creep in when you’re making modifications like that”. Similarly, P6 explained, “I really liked the ability to create relations between characters […] It’s something that made it a lot easier than any sort of other systems that I’ve used” and later added “[the tool] could be useful as a touch-up tool for when I want to make larger sweeping changes to a very long existing story […] I can confirm whether or not I want to keep each of these edits.”. This idea of reviewing the changes was important for participants, as P3 explained “You’d still have to review it, but it made enough changes that it [the story] was able to flow well.”
Visual writing can help explore and be creative
Many participants felt the tool was useful as a creative assistant. P6 explained “A lot of times, I am interested in how making a change to a given detail in a story would look […] I was appreciating it [the tool] as sort of an exploratory tool where even if it doesn’t end up producing something that I would personally publish, I think it’s really good to give an example that I can then engage with”. Similarly, P5 mentioned “You can just like, throw in a verb or the general action, interaction between them and it was kind of like spitballing ideas”. Regarding the timeline, P7 felt that “it’s an easy way to explore your options of what it sounds like if something happens here or if it sounds better when it’s there.”. They also mentioned how just looking at the visualization could help, for instance: “In terms of ideas, I could see how looking at the map, something could pop up into your head […] once you have, like, those objects laid out on a map, then that can help you spark inspiration”.
6.1. Observations and Participants’ Comments
6.1.1. Participants had different preferences
For instance, P4 mentioned “I really really like having a visual representation of the story, but this is not the visual representation I would choose […] I would like a combination of those two things [timeline and entity view] where those interlinks and connections appear within the timeline.”. Similarly, P6 explained how a single timeline would not be enough for them: “my brain interprets timelines in stories where I enjoy having things happening in multiple places at the same time and having a single linear approach makes it very hard to have that sort of interesting kind of development”. Other than visual representations, P4 also mentioned they “don’t want the system overwriting […] because then it’s not me expressing myself through text” and instead, they would prefer “just highlighting the places where things need to change, or the things are inconsistent”. This corroborates previous research on creative support tools needing to offer “many paths and many styles” to do tasks because users are unpredictable and have different preferences (Shneiderman et al., 2006).
6.1.2. Participants would have liked more control in using the visual representations to edit the story
While participants could express their intents with the tool, there are nuances in how an idea can be expressed in terms of writing style and words used, which is not currently supported by our system. For example, when discussing adding actions by connecting two entities, P8 mentioned “I feel like that’s very limiting how you wanted to express what you want versus just typing it out in a sentence because it kind of format it in a specific manner”. Similarly, P7 explained “in terms of expressing myself, I would say that it does help […] but I don’t think it allows for like total freedom of expression […] if you had something like, in your mind of how you wanted to say it, it might not come out that way”.
6.2. Limitations
6.2.1. The participants in our study were experienced and hesitant to using AI
All the participants in our study reported having multiple years of creative writing experience. Similarly, on a scale from 1-never to 5-always, participants rated their frequency of use of AI services a median of one, with five participants reporting never using AI services. P4 explained “I write because I want to, so if I want to be writing then there’s ideas that I have, and I want to express them. I don’t want a system to tell me what it is that I am going to be expressing”. This confirms previous findings that writers have different opinions towards AI (Hertzmann, 2018; Biermann et al., 2022) and those opinions might have impacted our results.
6.2.2. The system and the representations do not scale to long stories
One main limitation of the representations used in the system is that they quickly become too crowded and unusable for long stories. Currently, the system solves this problem through filtering mechanisms when the cursor hovers over a sentence or a scene. This results in hiding the elements irrelevant to the passage currently being selected. But this only works for local edits. When the edits need to happen between scenes far apart in the story, then it becomes challenging to find the right scene or the correct location. We believe that, for these cases, new aggregated representations and semantic zooming would be necessary.
6.2.3. The system might not have been robust enough to test the full potential of visual writing
While the system was optimized to be robust, it still depends on a large language model to understand the story and the instructions properly. In our tests, but also during the study, we observed the system sometimes responding unexpectedly by modifying elements it should not or by refusing to do certain modifications. Additionally, despite our best efforts to parallelize requests and speed up the execution, some manipulations were still slow to execute (up to 10-15 seconds at times). These issues might have confounded the results of our study. For example, it is known that to promote user exploration, interactions should be fast and with effects immediately visible (Masson et al., 2022; Shneiderman et al., 2006; Shneiderman, 1983).
6.3. Future Work
6.3.1. Visual writing to edit writing style and story plots
Creative writing is more than events, actors, locations, and time. We believe that visual writing could also help editing other aspects, such as the style of the writing. As proposed by Kempen et al. (1987), the syntax tree of the sentences could be made visible and be manipulable to restructure sentences. Alternatively, the tone, the length of the sentences, or the sentence structure could be represented using histograms or by embellishing the text. Then, manipulating these visuals would allow tweaking the style of the story. At a higher level, visual writing could also help revise the plot of the story. Similar to work on manipulating outlines (Dang et al., 2022a; Zhang et al., 2023), the idea would be to represent the main plot events of the story and then make them manipulable to reorder or change them.
6.3.2. Exploring new visual interactions
The idea of visual writing is broader than the representations and interactions proposed by our system. Any visual representation better suited for a task could be beneficial. For example, one could imagine using images to give the system an indication of what an object, character, or location look like. Then, the image could be modified to update the story (e.g., painting the hair of a character blond would edit the story so that the description matches the new hair colour). The system could also allow highlighting part of the image to indicate which aspect should be more or less emphasized in the description. Similarly, while our implementation uses LLM, visual writing could be implemented without and made more or less autonomous (Morris et al., 2024). Using LLM comes about with many challenges related to biases, privacy, and plagiarism (Weidinger et al., 2021; Meyer et al., 2023). An alternative, suggested by P4, could be to have the visual writing system only highlight the parts that should be modified to help the writer identify and edit them themselves.
6.3.3. Suggestions and creativity support in the visual space
The mental processes involved in interpreting visuals differ from those involved in processing language (Paivio, 1990). Future work could explore the impact on creativity of writing stories visually rather than with text. Similarly, traditional creativity support techniques for writers would be interesting to adapt to the visual space. For instance, there are many benefits of phrase suggestions (Buschek et al., 2021; Bhat et al., 2023), and future work could investigate if there are similar benefits to suggestions in the visual space. These visual suggestions would be shown when an interaction with a visual element is initiated to highlight different alternatives and to “auto-complete” the interaction.
7. Conclusion
We proposed visual writing, an approach to writing by manipulating visual representations to express editing intents. Through a taxonomy of possible visual representations based on narratology and a prototype implementation of some of these visual representations, we demonstrated one possible design for visual writing tools. A study with eight creative writers showed the potential of creative writing to help keep track of story elements, rapidly specify edits, and explore story variations in a way that encourages creativity. Broadly, our work advocates for a new generation of writing support tools that allow text editing through direct manipulation of visual representations.
Acknowledgements.
This work was supported by NSERC under grant RGPIN-2018-05072.References
- (1)
- Ackerman (2017) Angela Ackerman. 2017. The Efficient Writer: Using Timelines to Organize Story Details.
- Arnold et al. (2021) Kenneth C Arnold, April M Volzer, and Noah G Madrid. 2021. Generative Models Can Help Writers without Writing for Them. In IUI Workshops.
- Bal (1997) Mieke Bal. 1997. Narratology: Introduction to the Theory of Narrative (2nd edition ed.). University of Toronto Press, Scholarly Publishing Division, Toronto.
- Bhat et al. (2023) Advait Bhat, Saaket Agashe, Parth Oberoi, Niharika Mohile, Ravi Jangir, and Anirudha Joshi. 2023. Interacting with Next-Phrase Suggestions: How Suggestion Systems Aid and Influence the Cognitive Processes of Writing. In Proceedings of the 28th International Conference on Intelligent User Interfaces (IUI ’23). Association for Computing Machinery, New York, NY, USA, 436–452. https://doi.org/10.1145/3581641.3584060
- Biermann et al. (2022) Oloff C. Biermann, Ning F. Ma, and Dongwook Yoon. 2022. From Tool to Companion: Storywriters Want AI Writers to Respect Their Personal Values and Writing Strategies. In Designing Interactive Systems Conference. ACM, Virtual Event Australia, 1209–1227. https://doi.org/10.1145/3532106.3533506
- Brath (2021) Richard Brath. 2021. Surveying Wonderland for Many More Literature Visualization Techniques. https://arxiv.org/abs/2110.08584v1.
- Burnett and McIntyre (1995) Margaret M. Burnett and David W. McIntyre. 1995. Visual Programming. Computer 28, 03 (March 1995), 14–16. https://doi.org/10.1109/MC.1995.10027
- Buschek et al. (2021) Daniel Buschek, Martin Zürn, and Malin Eiband. 2021. The Impact of Multiple Parallel Phrase Suggestions on Email Input and Composition Behaviour of Native and Non-Native English Writers. In Proceedings of the 2021 CHI Conference on Human Factors in Computing Systems (CHI ’21). Association for Computing Machinery, New York, NY, USA, 1–13. https://doi.org/10.1145/3411764.3445372
- Calderwood et al. (2020) Alex Calderwood, Vivian Qiu, Katy Ilonka Gero, and Lydia B. Chilton. 2020. How Novelists Use Generative Language Models: An Exploratory User Study.. In HAI-GEN+ User2agent@ IUI.
- Carroll et al. (2009) Erin A. Carroll, Celine Latulipe, Richard Fung, and Michael Terry. 2009. Creativity factor evaluation: towards a standardized survey metric for creativity support. In Proceedings of the Seventh ACM Conference on Creativity and Cognition (Berkeley, California, USA) (C&C ’09). Association for Computing Machinery, New York, NY, USA, 127–136. https://doi.org/10.1145/1640233.1640255
- Chakrabarty et al. (2024) Tuhin Chakrabarty, Vishakh Padmakumar, Faeze Brahman, and Smaranda Muresan. 2024. Creativity Support in the Age of Large Language Models: An Empirical Study Involving Emerging Writers. arXiv:2309.12570 [cs]
- Cherry and Latulipe (2014) Erin Cherry and Celine Latulipe. 2014. Quantifying the Creativity Support of Digital Tools through the Creativity Support Index. ACM Trans. Comput.-Hum. Interact. 21, 4 (June 2014), 21:1–21:25. https://doi.org/10.1145/2617588
- Chung et al. (2022) John Joon Young Chung, Wooseok Kim, Kang Min Yoo, Hwaran Lee, Eytan Adar, and Minsuk Chang. 2022. TaleBrush: Sketching Stories with Generative Pretrained Language Models. In CHI Conference on Human Factors in Computing Systems. ACM, New Orleans LA USA, 1–19. https://doi.org/10.1145/3491102.3501819
- Comberg ([n. d.]) David Comberg. [n. d.]. Kurt Vonnegut on Shape of Stories. https://www.youtube.com/watch?v=oP3c1h8v2ZQ
- Dang et al. (2022a) Hai Dang, Karim Benharrak, Florian Lehmann, and Daniel Buschek. 2022a. Beyond Text Generation: Supporting Writers with Continuous Automatic Text Summaries. In Proceedings of the 35th Annual ACM Symposium on User Interface Software and Technology. ACM, Bend OR USA, 1–13. https://doi.org/10.1145/3526113.3545672
- Dang et al. (2023) Hai Dang, Sven Goller, Florian Lehmann, and Daniel Buschek. 2023. Choice Over Control: How Users Write with Large Language Models Using Diegetic and Non-Diegetic Prompting. In Proceedings of the 2023 CHI Conference on Human Factors in Computing Systems. 1–17. https://doi.org/10.1145/3544548.3580969 arXiv:2303.03199 [cs]
- Dang et al. (2022b) Hai Dang, Lukas Mecke, Florian Lehmann, Sven Goller, and Daniel Buschek. 2022b. How to Prompt? Opportunities and Challenges of Zero- and Few-Shot Learning for Human-AI Interaction in Creative Applications of Generative Models. https://arxiv.org/abs/2209.01390v1.
- Freytag (1895) Gustav Freytag. 1895. Technique of the drama: An exposition of dramatic composition and art. S. Griggs.
- Gobert and Beaudouin-Lafon (2023) Camille Gobert and Michel Beaudouin-Lafon. 2023. Lorgnette: Creating Malleable Code Projections. In Proceedings of the 36th Annual ACM Symposium on User Interface Software and Technology (UIST ’23). Association for Computing Machinery, New York, NY, USA, 1–16. https://doi.org/10.1145/3586183.3606817
- Hempel et al. (2019) Brian Hempel, Justin Lubin, and Ravi Chugh. 2019. Sketch-n-Sketch: Output-Directed Programming for SVG. In Proceedings of the 32nd Annual ACM Symposium on User Interface Software and Technology (UIST ’19). Association for Computing Machinery, New York, NY, USA, 281–292. https://doi.org/10.1145/3332165.3347925
- Hertzmann (2018) Aaron Hertzmann. 2018. Can Computers Create Art? https://arxiv.org/abs/1801.04486v6.
- Hoque et al. (2023) Md Naimul Hoque, Bhavya Ghai, Kari Kraus, and Niklas Elmqvist. 2023. Portrayal: Leveraging NLP and Visualization for Analyzing Fictional Characters. In Proceedings of the 2023 ACM Designing Interactive Systems Conference. ACM, Pittsburgh PA USA, 74–94. https://doi.org/10.1145/3563657.3596000
- Hutchins et al. (1985) Edwin L. Hutchins, James D. Hollan, and Donald A. Norman. 1985. Direct Manipulation Interfaces. Human–Computer Interaction 1, 4 (Dec. 1985), 311–338. https://doi.org/10.1207/s15327051hci0104_2
- Kempen et al. (1987) Gerard Kempen, G. Ambeek, Peter Desain, L. Konst, and Koenraad De Smedt. 1987. Author Environments: Fifth Generation Text Processors. (Jan. 1987).
- Kim et al. (2018) Nam Wook Kim, Benjamin Bach, Hyejin Im, Sasha Schriber, Markus Gross, and Hanspeter Pfister. 2018. Visualizing Nonlinear Narratives with Story Curves. IEEE Transactions on Visualization and Computer Graphics 24, 1 (Jan. 2018), 595–604. https://doi.org/10.1109/TVCG.2017.2744118
- Ko and Myers (2006) Amy J. Ko and Brad A. Myers. 2006. Barista: An Implementation Framework for Enabling New Tools, Interaction Techniques and Views in Code Editors. In Proceedings of the SIGCHI Conference on Human Factors in Computing Systems (CHI ’06). Association for Computing Machinery, New York, NY, USA, 387–396. https://doi.org/10.1145/1124772.1124831
- Lee et al. (2024) Mina Lee, Katy Ilonka Gero, John Joon Young Chung, Simon Buckingham Shum, Vipul Raheja, Hua Shen, Subhashini Venugopalan, Thiemo Wambsganss, David Zhou, Emad A. Alghamdi, Tal August, Avinash Bhat, Madiha Zahrah Choksi, Senjuti Dutta, Jin L. C. Guo, Md Naimul Hoque, Yewon Kim, Simon Knight, Seyed Parsa Neshaei, Agnia Sergeyuk, Antonette Shibani, Disha Shrivastava, Lila Shroff, Jessi Stark, Sarah Sterman, Sitong Wang, Antoine Bosselut, Daniel Buschek, Joseph Chee Chang, Sherol Chen, Max Kreminski, Joonsuk Park, Roy Pea, Eugenia H. Rho, Shannon Zejiang Shen, and Pao Siangliulue. 2024. A Design Space for Intelligent and Interactive Writing Assistants. https://doi.org/10.1145/3613904.3642697 arXiv:2403.14117 [cs]
- Maharjan et al. (2018) Suraj Maharjan, Sudipta Kar, Manuel Montes, Fabio A. González, and Thamar Solorio. 2018. Letting Emotions Flow: Success Prediction by Modeling the Flow of Emotions in Books. In Proceedings of the 2018 Conference of the North American Chapter of the Association for Computational Linguistics: Human Language Technologies, Volume 2 (Short Papers), Marilyn Walker, Heng Ji, and Amanda Stent (Eds.). Association for Computational Linguistics, New Orleans, Louisiana, 259–265. https://doi.org/10.18653/v1/N18-2042
- Marshall (2014) Colin Marshall. 2014. How J.K. Rowling Plotted Harry Potter with a Hand-Drawn Spreadsheet | Open Culture.
- Marti et al. (2018) Marcel Marti, Jodok Vieli, Wojciech Witoń, Rushit Sanghrajka, Daniel Inversini, Diana Wotruba, Isabel Simo, Sasha Schriber, Mubbasir Kapadia, and Markus Gross. 2018. CARDINAL: Computer Assisted Authoring of Movie Scripts. In Proceedings of the 23rd International Conference on Intelligent User Interfaces (IUI ’18). Association for Computing Machinery, New York, NY, USA, 509–519. https://doi.org/10.1145/3172944.3172972
- Masson et al. (2023) Damien Masson, Sylvain Malacria, Géry Casiez, and Daniel Vogel. 2023. Statslator: Interactive Translation of NHST and Estimation Statistics Reporting Styles in Scientific Documents. In Proceedings of the 36th Annual ACM Symposium on User Interface Software and Technology (UIST ’23). Association for Computing Machinery, New York, NY, USA, 1–14. https://doi.org/10.1145/3586183.3606762
- Masson et al. (2024) Damien Masson, Sylvain Malacria, Géry Casiez, and Daniel Vogel. 2024. DirectGPT: A Direct Manipulation Interface to Interact with Large Language Models. In Proceedings of the CHI Conference on Human Factors in Computing Systems (CHI ’24). Association for Computing Machinery, New York, NY, USA, 1–16. https://doi.org/10.1145/3613904.3642462
- Masson et al. (2022) Damien Masson, Jo Vermeulen, George Fitzmaurice, and Justin Matejka. 2022. Supercharging Trial-and-Error for Learning Complex Software Applications. In Proceedings of the 2022 CHI Conference on Human Factors in Computing Systems (CHI ’22). Association for Computing Machinery, New York, NY, USA, 1–13. https://doi.org/10.1145/3491102.3501895
- Mayer et al. (2018) Mikaël Mayer, Viktor Kuncak, and Ravi Chugh. 2018. Bidirectional Evaluation with Direct Manipulation. Implentation for: Bidirectional Evaluation with Direct Manipulation 2, OOPSLA (Oct. 2018), 127:1–127:28. https://doi.org/10.1145/3276497
- Meyer et al. (2023) Jesse G. Meyer, Ryan J. Urbanowicz, Patrick C. N. Martin, Karen O’Connor, Ruowang Li, Pei-Chen Peng, Tiffani J. Bright, Nicholas Tatonetti, Kyoung Jae Won, Graciela Gonzalez-Hernandez, and Jason H. Moore. 2023. ChatGPT and Large Language Models in Academia: Opportunities and Challenges. BioData Mining 16, 1 (July 2023), 20. https://doi.org/10.1186/s13040-023-00339-9
- Morris et al. (2024) Meredith Ringel Morris, Jascha Sohl-dickstein, Noah Fiedel, Tris Warkentin, Allan Dafoe, Aleksandra Faust, Clement Farabet, and Shane Legg. 2024. Levels of AGI for Operationalizing Progress on the Path to AGI. https://doi.org/10.48550/arXiv.2311.02462 arXiv:2311.02462 [cs]
- Munroe (2009) Randall Munroe. 2009. Movie Narrative Charts. https://xkcd.com/657/.
- Myers (1986) B. A. Myers. 1986. Visual Programming, Programming by Example, and Program Visualization: A Taxonomy. SIGCHI Bull. 17, 4 (April 1986), 59–66. https://doi.org/10.1145/22339.22349
- Neuwirth and Kaufer (1989) C. M. Neuwirth and D. S. Kaufer. 1989. The Role of External Representation in the Writing Process: Implications for the Design of Hypertext-Based Writing Tools. In Proceedings of the Second Annual ACM Conference on Hypertext - HYPERTEXT ’89. ACM Press, Pittsburgh, Pennsylvania, United States, 319–341. https://doi.org/10.1145/74224.74250
- NextUI (2021) nextui-org 2021. NextUI. nextui-org. https://nextui.org/
- OpenAI (2023) OpenAI. 2023. OpenAI Node API Library. OpenAI.
- Paivio (1990) Allan Paivio. 1990. Mental Representations: A Dual Coding Approach. Oxford University Press.
- Pope (1998) Thomas Pope. 1998. Good scripts, bad scripts: learning the craft of screenwriting through 25 of the best and worst films in history. Three Rivers Press.
- Qiang and Bingjie (2016) Lu Qiang and Chai Bingjie. 2016. StoryCake: A Hierarchical Plot Visualization Method for Storytelling in Polar Coordinates. In 2016 International Conference on Cyberworlds (CW). IEEE, Chongqing, China, 211–218. https://doi.org/10.1109/CW.2016.43
- React (2013) Meta 2013. React – A JavaScript Library for Building User Interfaces. Meta. https://reactjs.org/
- Reagan et al. (2016) Andrew J. Reagan, Lewis Mitchell, Dilan Kiley, Christopher M. Danforth, and Peter Sheridan Dodds. 2016. The Emotional Arcs of Stories Are Dominated by Six Basic Shapes. EPJ Data Science 5, 1 (Dec. 2016), 1–12. https://doi.org/10.1140/epjds/s13688-016-0093-1
- Reza et al. (2024) Mohi Reza, Nathan Laundry, Ilya Musabirov, Peter Dushniku, Zhi Yuan ”Michael” Yu, Kashish Mittal, Tovi Grossman, Michael Liut, Anastasia Kuzminykh, and Joseph Jay Williams. 2024. ABScribe: Rapid Exploration & Organization of Multiple Writing Variations in Human-AI Co-Writing Tasks Using Large Language Models. https://doi.org/10.1145/3613904.3641899 arXiv:2310.00117 [cs]
- SaveTheCat (2014) Blake Snyder Enterprises 2014. Save The Cat! Story Structure Software. Blake Snyder Enterprises. https://savethecat.com/save-the-cat-software
- Shneiderman (1983) Shneiderman. 1983. Direct Manipulation: A Step Beyond Programming Languages. Computer 16, 8 (Aug. 1983), 57–69. https://doi.org/10.1109/MC.1983.1654471
- Shneiderman (1982) Ben Shneiderman. 1982. The Future of Interactive Systems and the Emergence of Direct Manipulation†. Behaviour & Information Technology 1, 3 (July 1982), 237–256. https://doi.org/10.1080/01449298208914450
- Shneiderman et al. (2006) Ben Shneiderman, Gerhard Fischer, Mary Czerwinski, Mitch Resnick, Brad Myers, Linda Candy, Ernest Edmonds, Mike Eisenberg, Elisa Giaccardi, Tom Hewett, Pamela Jennings, Bill Kules, Kumiyo Nakakoji, Jay Nunamaker, Randy Pausch, Ted Selker, Elisabeth Sylvan, and Michael Terry. 2006. Creativity Support Tools: Report From a U.S. National Science Foundation Sponsored Workshop. International Journal of Human–Computer Interaction 20, 2 (May 2006), 61–77. https://doi.org/10.1207/s15327590ijhc2002_1
- SlateJS (2016) Slate.js 2016. Slate.js. Slate.js. https://www.slatejs.org
- Suh et al. (2024) Sangho Suh, Meng Chen, Bryan Min, Toby Jia-Jun Li, and Haijun Xia. 2024. Luminate: Structured Generation and Exploration of Design Space with Large Language Models for Human-AI Co-Creation. In Proceedings of the CHI Conference on Human Factors in Computing Systems (Honolulu, HI, USA) (CHI ’24). Association for Computing Machinery, New York, NY, USA, Article 644, 26 pages. https://doi.org/10.1145/3613904.3642400
- Tanahashi and Kwan-Liu Ma (2012) Y. Tanahashi and Kwan-Liu Ma. 2012. Design Considerations for Optimizing Storyline Visualizations. IEEE Transactions on Visualization and Computer Graphics 18, 12 (Dec. 2012), 2679–2688. https://doi.org/10.1109/TVCG.2012.212
- Watson et al. (2019) Katie Watson, Samuel S. Sohn, Sasha Schriber, Markus Gross, Carlos Manuel Muniz, and Mubbasir Kapadia. 2019. StoryPrint: An Interactive Visualization of Stories. In Proceedings of the 24th International Conference on Intelligent User Interfaces (IUI ’19). Association for Computing Machinery, New York, NY, USA, 303–311. https://doi.org/10.1145/3301275.3302302
- Weidinger et al. (2021) Laura Weidinger, John Mellor, Maribeth Rauh, Conor Griffin, Jonathan Uesato, Po-Sen Huang, Myra Cheng, Mia Glaese, Borja Balle, Atoosa Kasirzadeh, Zac Kenton, Sasha Brown, Will Hawkins, Tom Stepleton, Courtney Biles, Abeba Birhane, Julia Haas, Laura Rimell, Lisa Anne Hendricks, William Isaac, Sean Legassick, Geoffrey Irving, and Iason Gabriel. 2021. Ethical and Social Risks of Harm from Language Models. https://doi.org/10.48550/arXiv.2112.04359 arXiv:2112.04359 [cs]
- Williams ([n. d.]) Eric Williams. [n. d.]. How to View and Appreciate Great Movies. https://www.thegreatcourses.com/courses/how-to-view-and-appreciate-great-movies
- Yuan et al. (2022) Ann Yuan, Andy Coenen, Emily Reif, and Daphne Ippolito. 2022. Wordcraft: Story Writing With Large Language Models. In Proceedings of the 27th International Conference on Intelligent User Interfaces (IUI ’22). Association for Computing Machinery, New York, NY, USA, 841–852. https://doi.org/10.1145/3490099.3511105
- Zhang et al. (2023) Zheng Zhang, Jie Gao, Ranjodh Singh Dhaliwal, and Toby Jia-Jun Li. 2023. VISAR: A Human-AI Argumentative Writing Assistant with Visual Programming and Rapid Draft Prototyping. In Proceedings of the 36th Annual ACM Symposium on User Interface Software and Technology. ACM, San Francisco CA USA, 1–30. https://doi.org/10.1145/3586183.3606800
- Zhu (2013) Xiaojin Zhu. 2013. Persistent Homology: An Introduction and a New Text Representation for Natural Language Processing. In IJCAI International Joint Conference on Artificial Intelligence. 1953–1959.
- Zhu and Kolassa (2018) Yaqian Zhu and John Kolassa. 2018. Assessing and Comparing the Accuracy of Various Bootstrap Methods. Communications in Statistics - Simulation and Computation 47, 8 (Sept. 2018), 2436–2453. https://doi.org/10.1080/03610918.2017.1348516